OpenStudy (vshiroky):

Write an equation for the line passing through the given points. Give the final answer A)slope-intercept form and B) Standard form (-3,-1) and (4,-6)

OpenStudy (vshiroky):

I had to start over. I'm so lost!

OpenStudy (anonymous):

Do you know that forms?

OpenStudy (skullpatrol):

Go To YouTube and look at Khan Academy slope intercept form of a line

OpenStudy (vshiroky):

I did -6+3/2+5=3/7

OpenStudy (skullpatrol):

https://www.youtube.com/watch?v=uk7gS3cZVp4

OpenStudy (anonymous):

Slope-intercept is y=mx+b .....

OpenStudy (anonymous):

You have to find the slope... Use the equation y2-y1/x2-x1 .... Plug those in and get a slope

OpenStudy (jdoe0001):

\(\bf \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ &({\color{red}{ -3}}\quad ,&{\color{blue}{ -1}})\quad &({\color{red}{ 4}}\quad ,&{\color{blue}{ -6}}) \end{array} \\\quad \\ slope = {\color{green}{ m}}= \cfrac{rise}{run} \implies \cfrac{{\color{blue}{ y_2}}-{\color{blue}{ y_1}}}{{\color{red}{ x_2}}-{\color{red}{ x_1}}} \\ \quad \\ y-{\color{blue}{ y_1}}={\color{green}{ m}}(x-{\color{red}{ x_1}})\qquad \textit{plug in the values and solve for "y"}\\ \qquad \uparrow\\ \textit{point-slope form} \)

OpenStudy (skullpatrol):

https://www.youtube.com/watch?v=-6Fu2T_RSGM @vshiroky

OpenStudy (anonymous):

Then plug it into point-slope... y-y1=m(x-x1) .... Then solve

OpenStudy (vshiroky):

Right, but I keep getting a decimal

OpenStudy (vshiroky):

It's not coming out right.

OpenStudy (jdoe0001):

leave the slope as a fraction, and use it like so

OpenStudy (vshiroky):

Right, but when I type it in it's not right.

OpenStudy (vshiroky):

I got m=5/7

OpenStudy (jdoe0001):

hmm what have you gotten thus far?

OpenStudy (vshiroky):

y+1=5/7(x+3)

OpenStudy (vshiroky):

When I do 5/7(3) it's coming up as a decimal

OpenStudy (jdoe0001):

well, close \(\bf slope = {\color{green}{ m}}= \cfrac{rise}{run} \implies \cfrac{{\color{blue}{ -6}}-{\color{blue}{ -1}}}{{\color{red}{ 4}}-{\color{red}{ -3}}}\implies \cfrac{-5}{7}\)

OpenStudy (jdoe0001):

recall the "6" is negative

OpenStudy (vshiroky):

Ok, but I'm still getting a decimal when I do -5/7(3)

OpenStudy (jdoe0001):

well, when you distribute the slope, leave the product as fraction

OpenStudy (skullpatrol):

where did the (3) come from?

OpenStudy (jdoe0001):

you'd get decimal if you divide them, yes

OpenStudy (vshiroky):

Because I did y+1=-5/7(x+3)

OpenStudy (skullpatrol):

So now multiply out y+1=(-5/7)(x+3)

OpenStudy (vshiroky):

What I don't get is when i do -5/7(3) I get a decimal

OpenStudy (vshiroky):

I keep getting -2.14

OpenStudy (skullpatrol):

Just leave it as a fraction

OpenStudy (vshiroky):

I don't know how to do that when I am doing it from my phone. It doesn't do fractions

OpenStudy (vshiroky):

so what y=-5/7x-15/7?

OpenStudy (skullpatrol):